OpenStudy (anonymous):

\[ \frac{ 15^{2} * 8^{-1} }{ 6^{3} * 10^{2}}\]

OpenStudy (anonymous):

without calculator

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

make 15*15 divide by 8 than 15*15 multiply by 8 to have same base

OpenStudy (anonymous):

i dont understand

hartnn (hartnn):

write 15 = 3*5 6 = 2*3 and use \(\large (xy)^m=x^my^m\)

OpenStudy (anonymous):

okk